A recent study by the World Bank wished to determine whether there was a relationship between the abundance of natural resources in a country and its long term rate of economic growth. Their study used 40 (n) countries over a long period to 1995. Letting Y be the rate of growth measured as a percentage and X be a measure of natural resource abundance, the following relationship between the two variables was estimated. Standard errors of b0 and b1 are reported in parentheses under the coefficients for b0 and b1 Yhat = 5.664 - 0.0059 X (1.285) (0.0029) To test the significance of the relationship, a 5% level of significance was adopted. 1. State the direction of the alternative hypothesis used to test the statistical significance of the relationship. Type gt (greater than), ge (greater than or equal to), lt (less than), le (less than or equal to) or ne (not equal to) as appropriate in the box. 2. Use the tables in the textbook to determine the critical value of the test statistic (in absolute terms) to three decimal places 3. Calculate the test statistic, reporting your answer to two decimal places. 4. Is the null hypothesis rejected for this test? Type yes or no. 5. Disregarding your answer in part 4, if the null hypothesis was rejected, would there appear to be a linear relationship between resource abundance and economic growth? Type yes or no
Answer #1

Solution:

1) direction of the alternative hypothesis used to test the statistical significance of the relationship.

H0: The slope of the regression line is equal to zero.
    Ha: The slope of the regression line is not equal to zero

Ans) ne

2) the critical value of the test statistic

We get the slope (b1) and the standard error (SE) from the regression output.

b1 = -0.0059 SE = 0.0029

Test stat t = b1/SE = -0.0059 /0.0029 = -2.03

3) Calculate the test statistic

Degrees of Freedom = n - 2 = 40 - 2 = 38

   critical value (in absolute terms) for alpha = 0.05 and df = 38 is 2.024

4) Is the null hypothesis rejected for this test

The rejection area is t < - 2.024 or t > +2.024

Since test stat does not lie in the rejection region we do not reject the Null

Ans) No

5) there appear to be a linear relationship between resource abundance and economic growth

No